# Practice Makes Perfect: Planning to Learn Skill Parameter Policies

### v1

@@ Abstract
One promising approach towards effective robot decision making in complex, long-horizon tasks is to sequence together parameterized skills. We consider a setting where a robot is initially equipped with (1) a library of parameterized skills, (2) an AI planner for sequencing together the skills given a goal, and (3) a very general prior distribution for selecting skill parameters. Once deployed, the robot should rapidly and autonomously learn to improve its performance by specializing its skill parameter selection policy to the particular objects, goals, and constraints in its environment. In this work, we focus on the active learning problem of choosing which skills to practice to maximize expected future task success. We propose that the robot should estimate the competence of each skill, extrapolate the competence (asking: “how much would the competence improve through practice?”), and situate the skill in the task distribution through competence-aware planning. This approach is implemented within a fully autonomous system where the robot repeatedly plans, practices, and learns without any environment resets. Through experiments in simulation, we find that our approach learns effective parameter policies more sample-efficiently than several baselines. Experiments in the real-world demonstrate our approach’s ability to handle noise from perception and control and improve the robot’s ability to solve two long-horizon mobile-manipulation tasks after a few hours of autonomous practice. Project website: https://planning-to-practice.github.io $*{}^{*}$qual contribution. Work done during internship at The AI Institute. Correspondence to {njk, tslvr}@csail.mit.edu.
@@ II-A Modelling the World
The action space of the MDP is defined by a set of parameterized skills that have continuous parameters. It is often convenient to define object parameters as well, but for the purpose of this work, these play a minor role, so we treat them as part of the skill except where otherwise noted. For example, in the Ball-Ring environment, Place($\circ$) is one skill $u\in\mathcal{U}$, where $\circ$ denotes a placeholder for a continuous parameter, and Place($\circ$) is another $u^{\prime}\in\mathcal{U}$. The continuous parameters for both skills are xy relative offsets between the gripper and target surface (the height and gripper orientation are fixed).
@@ II-B Planning to Solve Tasks
Given a task, the robot will plan to generate actions that are likely to accomplish the goal from the initial state. Following previous work [55, 1, 51, 35], we decompose planning into two levels: skill sequencing and continuous parameter selection. Skill sequencing consists of generating a skeleton, e.g., ((((MoveTo($\circ$), Pick(ball, $\circ$), MoveTo($\circ$), Place(ball, $\circ$))normal-))). Given a skeleton, we select continuous parameters using parameter policies $\theta\sim\pi_{u}(\cdot\mid x)$. Since parameter selection is conditioned on the state $x$, and since we are not assuming a known transition distribution [51], we sample and execute each skill greedily. If the skill terminates and does not meet its success condition, we replan. See Algorithm 1 for a summary.
@@
The competence $c_{u,\pi}$ of a skill $u$ with current parameter policy $\pi_{u}$ is the expected success $\mathop{\mathbb{E}}[J_{u}(X_{t+1})\mid I_{u}(X_{t})]$, where $X_{t}$ is a random variable for the state before skill execution, $A_{t}$ is a r.v. for the action generated from $\theta\sim\pi_{u}(\cdot\mid X_{t})$, and $X_{t+1}\sim P(\cdot\mid X_{t},A_{t})$.
@@
654321 Input: Current state $x$ and goal $g$. Generate a skeleton $(u_{0},\dots,u_{n})$ to $g$ from $x$. For $i=0,\dots,n$:  Sample $\theta\sim\pi_{u_{i}}$ and execute $u_{i}(\theta)$.  Perceive and update the current state $x$.  If $J_{u_{i}}(x)\neq 1$, repeat from line 2 (replan).
@@
7654321 Initialize parameter policies $\Pi=\{\pi^{0}_{u}:u\in\mathcal{U}\}$. Repeat:  If a human has given a goal $g$:   Plan and execute to $g$ with Algorithm 1.  Else:   Select and execute actions of the robot’s choice.  Update $\Pi$ every $m$ iterations.
@@ III Planning to Learn
7654321 Input: Current parameter policies $\Pi$. Select a skill $u\in\mathcal{U}$ to practice (see Algorithm 4). Plan to $I_{u}$ using Algorithm 1 with $\Pi$. Practice the skill $u$ one time:  Sample parameters $\theta$ from an explore policy $\pi^{+}_{u}$.  Execute $u(\theta)$ and record the transition. Repeat from line 2 until free time expires.
@@ III-A Selecting Skills to Practice
In other words, the expected probability that planning succeeds without replanning over the task distribution.
@@
321 For each $u\in\mathcal{U}$ with current parameter policy $\pi_{u}$:  Estimate the current competence $c_{u,\pi}$.  Extrapolate: predict $c_{u,\pi^{\prime}}$, the competence after
4  practicing $u$ and updating $\pi_{u}$ to $\pi_{u}^{\prime}$.  Situate the competence in the task distribution,
5  computing $J_{\text{skill}}(u)\triangleq J_{\text{tasks}}(\Pi-\{\pi\}\cup\{\pi^{\prime}\})$. Return $\operatorname*{argmax}_{u}J_{\text{skill}}(u)$ for practice.
@@ IV Experiments
Figure 3: Simulation results. Percentage of evaluation tasks solved vs. number of online transitions collected for all approaches in all simulated environments. Solid lines represent means and shading represents standard error across 10 seeds.
@@
Results and Analysis. Figure 3 shows a plot of the evaluation success rate of all approaches in all simulated environments. EES is consistently the most sample efficient, achieving higher success rates after fewer online transitions than the baselines. Fail Focus falters not only because it focuses on impossible skills, like jumping in Light Switch or placing the ball on the table in Ball Ring, but also because it is not sufficiently situated in the task distribution. For example, there are no impossible skills in Cleanup Playroom, but EES still outperforms Fail Focus. Competence Gradient is competitive with Fail Focus but similarly lacks situatedness. The poor performance of Skill Diversity, Task-Relevant, Task Repeat, and Random Skills underscore the importance of directed active practice in these long-horizon environments with many possible skills. Like the Random Skills baseline, MAPLE-Q fails to solve any evaluation tasks, which is not surprising given the highly limited number of online transitions that we are considering; previous related work [44] learns from multiple orders of magnitude more data. We verified that our implementation does well in far simpler environments with far more data; see Appendix -E for further discussion.
@@
Tables I and II show EES evaluation results for the real-robot environments. Our approach is able to improve its performance after 120 and 240 real-world skill-executions respectively. Each seed took 1-3 hours of real robot time; see the supplementary video. These results are especially noteworthy given the non-trivial noise in perception and control, leading to both false positives and false negatives in the skill datasets ($\mathcal{D}_{u}$) collected for parameter policy learning.
@@ V-A Exploration in Reinforcement Learning
The problem of sequentially selecting actions that lead to efficient learning is central to exploration in reinforcement learning [29, 46, 13, 8, 2]. One important difference between our setting and RL is that we do not have a temporal credit assignment problem: given skill success conditions and Assumption 1, each parameter policy learning problem is self-contained. Our skill selection problem is therefore related to exploration in multi-armed bandits [9], but different still, since selecting a skill to practice does not lead to a task reward, but rather, to a data point that can be used to improve the parameter policy for that skill. Note that the inner problem of choosing parameters for a selected skill is a bandit problem (with infinite arms) [10] as explained in Section III-B, but our primary interest is the outer skill selection problem.
@@ VI Limitations and Future Work
There are several limitations of the present work and of EES as a general method. For the sake of rapidly learning on a real robot, we started with a considerable amount of prior knowledge: known object (feature) detectors, fully-specified parameterized skills (and operators for planning), low-dimensional feature selectors for parameter policy training, and good parameter priors. Previous work has considered learning each of these components (e.g.,  [52, 54]), but doing so may require significantly more data than what we considered here. We also made Assumptions 1-3, which are strong, and while we need not satisfy them completely to attain good performance, they remain worthy of further scrutiny. Our relatively naive treatment of noise and our assumption of full observability are also clearly limiting. Furthermore, our commitment to planning to practice (Algorithm 3) is perhaps overly myopic: better strategies might anticipate that practicing one skill enables quickly practicing another, reasoning over sequences of practice attempts. Finally, this work presupposes that robots should be practicing and learning during deployment. The extent to which this is true depends greatly on the nature of the deployment and the constraints under which the robot is allowed to practice.
One future direction that could address multiple limitations simultaneously would be to give the agent access to a simulator. Even if the simulator were a coarse approximation of the real world, the robot could nonetheless use it to bootstrap real-world practice time. The same simulator could be used for reasoning about potentially irreversible actions before executing them in the real world, and for integrated task and motion planning (TAMP) [24] towards removing Assumption 1. Leveraging TAMP would also be a step toward more principled planning in stochastic [50] and partially-observable environments [23].
@@ Acknowledgments
We gratefully acknowledge support from NSF grant 2214177; from AFOSR grant FA9550-22-1-0249; from ONR MURI grant N00014-22-1-2740; from ARO grant W911NF-23-1-0034 and from the MIT Quest for Intelligence. Nishanth, Tom, and Willie are supported by NSF GRFP fellowships. We thank Will Shen for feedback and suggestions on an early paper draft, as well as invaluable help with creating the accompanying website. We also thank Russell Mendonca for helpful early discussions, especially with respect to the MAPLE-Q baseline. We are grateful for helpful discussion, brainstorming, and support from Stefanie Tellex, Ashay Athalye, Tushar Kusnur, Jiuguang Wang, Gustavo Goretkin, Andrew Messing, Joe St. Germain and others at the AI Institute. We thank Hannah Blumberg for invaluable early help with designing one of our robot domains, as well as helpful comments on an early draft of this paper. We acknowledge the MIT SuperCloud and Lincoln Laboratory Supercomputing Center for providing HPC resources that have contributed to the simulation results reported within this paper. Finally, we wish to thank our three Spot robots, Moana, Donner, and Kepler, for being so reliable throughout the extensive prototyping and experimentation required for this paper. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of our sponsors.

### v2

@@ Abstract
One promising approach towards effective robot decision making in complex, long-horizon tasks is to sequence together parameterized skills. We consider a setting where a robot is initially equipped with (1) a library of parameterized skills, (2) an AI planner for sequencing together the skills given a goal, and (3) a very general prior distribution for selecting skill parameters. Once deployed, the robot should rapidly and autonomously learn to improve its performance by specializing its skill parameter selection policy to the particular objects, goals, and constraints in its environment. In this work, we focus on the active learning problem of choosing which skills to practice to maximize expected future task success. We propose that the robot should estimate the competence of each skill, extrapolate the competence (asking: “how much would the competence improve through practice?”), and situate the skill in the task distribution through competence-aware planning. This approach is implemented within a fully autonomous system where the robot repeatedly plans, practices, and learns without any environment resets. Through experiments in simulation, we find that our approach learns effective parameter policies more sample-efficiently than several baselines. Experiments in the real-world demonstrate our approach’s ability to handle noise from perception and control and improve the robot’s ability to solve two long-horizon mobile-manipulation tasks after a few hours of autonomous practice. Project website: http://ees.csail.mit.edu Equal contribution. Work done during internship at The AI Institute. Correspondence to {njk, tslvr}@csail.mit.edu.
@@ II-A Modelling the World
The action space of the MDP is defined by a set of parameterized skills that have continuous parameters. It is often convenient to define object parameters as well, but for the purpose of simplifying exposition, we will treat these as part of the skill unless otherwise noted. For example, in the Ball-Ring environment, Place($\circ$) is one skill $u\in\mathcal{U}$, where $\circ$ denotes a placeholder for a continuous parameter, and Place($\circ$) is another $u^{\prime}\in\mathcal{U}$ [^1] Note that in implementation, we implement object-parameterized skills as discussed in Appendix Section -F. The continuous parameters for both skills are xy relative offsets between the gripper and target surface (the height and gripper orientation are fixed).
@@ II-B Planning to Solve Tasks
Given a task, the robot will plan to generate actions that are likely to accomplish the goal from the initial state. Following previous work [56, 1, 52, 35], we decompose planning into two levels: skill sequencing and continuous parameter selection. Skill sequencing consists of generating a skeleton, e.g., ((((MoveTo($\circ$), Pick(ball, $\circ$), MoveTo($\circ$), Place(ball, $\circ$))))). Given a skeleton, we select continuous parameters using parameter policies $\theta\sim\pi_{u}(\cdot\mid x)$. Since parameter selection is conditioned on the state $x$, and since we are not assuming a known transition distribution [52], we sample and execute each skill greedily. If the skill terminates and does not meet its success condition, we replan. See Algorithm 1 for a summary.
@@
The competence $c_{u,\pi}$ of a skill $u$ with current parameter policy $\pi_{u}$ is the expected success $\mathop{\mathbb{E}}[J_{u}(X_{t+1})\mid I_{u}(X_{t})]$, where $I_{u}$ characterizes the skill $u$ can be initiated from, $X_{t}$ is a random variable for the state before skill execution, $A_{t}$ is a r.v. for the action generated from $\theta\sim\pi_{u}(\cdot\mid X_{t})$, and $X_{t+1}\sim P(\cdot\mid X_{t},A_{t})$.
@@
1 Input: Current state $x$ and goal $g$.
2 Generate a skeleton $(u_{0},\dots,u_{n})$ to $g$ from $x$.
3 For $i=0,\dots,n$:
4  Sample $\theta\sim\pi_{u_{i}}$ and execute $u_{i}(\theta)$.
5  Perceive and update the current state $x$.
6  If $J_{u_{i}}(x)\neq 1$, repeat from line 2 (replan).
@@ 3 For $i=0,\dots,n$:
1 Initialize parameter policies $\Pi=\{\pi^{0}_{u}:u\in\mathcal{U}\}$.
2 Repeat:
3If a human has given a goal $g$:
4   Plan and execute to $g$ with Algorithm 1.
5Else:
6   Select and execute actions of the robot’s choice.
7  Update $\Pi$ every $m$ iterations.
@@ III Planning to Learn
1 Input: Current parameter policies $\Pi$.
2 Select a skill $u\in\mathcal{U}$ to practice (see Algorithm 4).
3 Plan to $I_{u}$ using Algorithm 1 with $\Pi$.
4 Practice the skill $u$ one time:
5  Sample parameters $\theta$ from an explore policy $\pi^{+}_{u}$.
6Execute $u(\theta)$ and record the transition.
7 Repeat from line 2 until free time expires.
@@ III-A Selecting Skills to Practice
that is, the expected probability that planning succeeds without replanning over the task distribution.
@@
1 For each $u\in\mathcal{U}$ with current parameter policy $\pi_{u}$:
2Estimate the current competence $c_{u,\pi}$.
3Extrapolate: predict $c_{u,\pi^{\prime}}$, the competence after
 practicing $u$ and updating $\pi_{u}$ to $\pi_{u}^{\prime}$.
4Situate the competence in the task distribution,
 computing $J_{\text{skill}}(u)\triangleq J_{\text{tasks}}(\Pi-\{\pi\}\cup\{\pi^{\prime}\})$.
5 Return $\operatorname*{argmax}_{u}J_{\text{skill}}(u)$ for practice.
@@ IV Experiments
Figure 3: Simulation results. Percentage of evaluation tasks solved vs. number of online transitions collected for all approaches in all simulated environments. Solid lines represent means and shading represents standard error across 10 seeds. Note that all approaches used the same parameter priors and feature engineering, discussed in detail in Appendix Sections -F and -C. We run additional ablation experiments on these choices in Appendix Section -G.
@@
Results and Analysis. Figure 3 shows a plot of the evaluation success rate of all approaches in all simulated environments. EES is consistently the most sample efficient, achieving higher success rates after fewer online transitions than the baselines. This is particularly evident when the number of online transitions is high enough for EES to practice useful skills a sufficient number of times for a noticeable improvement. Fail Focus falters not only because it focuses on impossible skills, like jumping in Light Switch or placing the ball on the table in Ball Ring, but also because it is not sufficiently situated in the task distribution. For example, there are no impossible skills in Cleanup Playroom, but EES still outperforms Fail Focus. Competence Gradient is competitive with Fail Focus but similarly lacks situatedness. The poor performance of Skill Diversity, Task-Relevant, Task Repeat, and Random Skills underscore the importance of directed active practice in these long-horizon environments with many possible skills. Like the Random Skills baseline, MAPLE-Q fails to solve any evaluation tasks, which is not surprising given the highly limited number of online transitions that we are considering; previous related work [45] learns from multiple orders of magnitude more data. We verified that our implementation does well in far simpler environments with far more data; see Appendix -E for further discussion.
Tables I and II show EES evaluation results for the real-robot environments. Our approach is able to improve its performance after 120 and 240 real-world skill-executions respectively. Each seed took 1-3 hours of real robot time; see the supplementary video. These results are especially noteworthy given the complex skills (e.g. sweeping from the table), relatively long-horizons necessary for both setting up practice and for solving each task, and non-trivial noise in perception and control, which lead to both false positives and false negatives in the skill datasets ($\mathcal{D}_{u}$) collected for parameter policy learning.
@@ V-A Exploration in Reinforcement Learning
The problem of sequentially selecting actions that lead to efficient learning is central to exploration in reinforcement learning [29, 47, 13, 8, 2]. One important difference between our setting and RL is that we do not have a temporal credit assignment problem: given skill success conditions and Assumption 1, each parameter policy learning problem is self-contained. Our skill selection problem is therefore related to exploration in multi-armed bandits [9], but different still, since selecting a skill to practice does not lead to a task reward, but rather, to a data point that can be used to improve the parameter policy for that skill. Note that the inner problem of choosing parameters for a selected skill is a bandit problem (with infinite arms) [10] as explained in Section III-B, but our primary interest is the outer skill selection problem. Our method for skill selection can also be viewed as a restricted form of curriculum learning for RL [44].
@@
Recent work by Vats et al. [60] considers a variation of estimation, extrapolation, and situation in the context of recovery learning, where additional skills are learned to supplement a given set of imperfect parameterized skills. Instead of asking which parameter policy to practice, they consider which recovery skill to practice. They introduce a “value of failures” objective that is similar to our $J_{\text{tasks}}$ objective. One important difference is that they assume access to a simulator; they therefore do not need to plan to practice.
@@ VI Limitations and Future Work
There are several limitations of the present work and of EES as a general method. For the sake of rapidly learning on a real robot, we started with a considerable amount of prior knowledge: known object (feature) detectors, fully-specified parameterized skills (and operators for planning), low-dimensional feature selectors for parameter policy training, and good parameter priors. Previous work has considered learning each of these components (e.g.,  [53, 55]), but doing so may require significantly more data than what we considered here. We also made Assumptions 1-3, which are strong, and while we need not satisfy them completely to attain good performance, they remain worthy of further scrutiny. Our approach also implicitly assumes the overall task distribution is stationary, which may not hold in many complex and dynamic real-world environments. Our relatively naive treatment of noise and our assumption of full observability are also clearly limiting. Furthermore, our commitment to planning to practice (Algorithm 3) is perhaps overly myopic: better strategies might anticipate that practicing one skill enables quickly practicing another, reasoning over sequences of practice attempts. Finally, this work presupposes that robots should be practicing and learning during deployment. The extent to which this is true depends greatly on the nature of the deployment and the constraints under which the robot is allowed to practice.
One future direction that could address multiple limitations simultaneously would be to give the agent access to a simulator. A significant challenge with doing this is that the precise models of all the objects the robot might encounter during deployment are not available ahead of time, and thus the robot must acquire aspects of these online. However, even if the acquired simulation were a coarse approximation of the real world, the robot could nonetheless use it to bootstrap real-world practice time. The same simulation could be used for reasoning about potentially irreversible actions before executing them in the real world, and for integrated task and motion planning (TAMP) [24] towards removing Assumption 1. Leveraging TAMP would also be a step toward more principled planning in stochastic [51] and partially-observable environments [23].
@@ Acknowledgments
We gratefully acknowledge support from NSF grant 2214177; from AFOSR grant FA9550-22-1-0249; from ONR MURI grant N00014-22-1-2740; from ARO grant W911NF-23-1-0034 and from the MIT Quest for Intelligence. Nishanth, Tom, and Willie are supported by NSF GRFP fellowships. We thank Will Shen for feedback and suggestions on an early paper draft, as well as invaluable help with creating the accompanying website. We also thank Russell Mendonca for helpful early discussions, especially with respect to the MAPLE-Q baseline. We are grateful for helpful discussion, brainstorming, and support from Stefanie Tellex, Ashay Athalye, Tushar Kusnur, Jiuguang Wang, Gustavo Goretkin, Andrew Messing, Joe St. Germain and others at the AI Institute. We thank Hannah Blumberg for invaluable early help with prototyping one of our robot domains, as well as helpful comments on an early draft of this paper. We also thank Chris Agia, Shivam Vats, and Peter Stone for helpful comments on an earlier version of the paper. We gratefully acknowledge the MIT SuperCloud and Lincoln Laboratory Supercomputing Center for providing HPC resources that have contributed to the simulation results reported within this paper. Finally, we wish to thank our three Spot robots, Moana, Donner, and Kepler, for being so reliable throughout the extensive prototyping and experimentation required for this paper. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of our sponsors.